# references.bib

### v2

@@ 4.2 Mixed Graph-level Tasks
4.3 Node-level Benchmarks
@@
Node-level tasks are an interesting challenge for our proposed approach as the PMA module is not needed and node-level propagation is the most natural strategy of learning node representations. To this end, we did prototype with an edge-to-node pooling module which would allow ESA to learn node embeddings, with good results. However, the approach does not currently scale to the millions of edges that some heterophilous graphs have. For these reasons, we revert to the simpler node-set attention (NSA) formulation. Table 7 summarizes our results, indicating that NSA performs well on all $11$ node-level benchmarks, including homophilous (citation), heterophilous, and shortest path tasks. We have adapted Graphormer and TokenGT for node classification as this functionality was not originally available, although they require integer node and edge features which restricts their use on some datasets (denoted by n/a in Table 7). NSA achieves the highest MCC score on homophilous and heterophilous graphs, but GCN has the highest accuracy on cora (Table 9). Some methods, for instance GraphSAGE [10.5555/3294771.3294869] are designed to perform particularly well under heterophily and to account for that we include GraphSAGE [10.5555/3294771.3294869] and Graph Transformer [ijcai2021p214] as the two top performing baselines from Platonov et al. [platonov2023a] in our extended results in Table 10. With the exception of the amazon ratings datasets, we outperform the two baselines by a noticeable margin. Our results on chameleon stand out in particular, as well as on the shortest path benchmarks, where other graph transformers are unable to learn and even PNA fails to be competitive.
4.4 Effects of Varying the Layer Order and Type
@@
In Table 8, we summarise the results of an ablation relative to the interleaving operator with different order and types of layers in our architecture. Smaller datasets perform well with $4$ to $6$ feature extraction layers, while larger datasets with more complex graphs like mnist benefit from up to $10$ layers. We have generally observed that the top configurations tend to include self-attention layers at the front, with masked attention layers in the middle and self-attention layers at the end, surrounding the PMA readout. Naive configurations such as all-masked layers or simply alternating masked and self-attention layers do not tend to be optimal for graph-level prediction tasks. This ablation experiment demonstrates the importance of vertical combination of masked and self-attention layers for the performance of our model.
4.5 Time and Memory Scaling
The theoretical properties of ESA are discussed in Methods, Section 3.3, where we also cover deep learning library limitations and possible optimisations (also see SI 5). Here, we have empirically evaluated the time and memory scaling of ESA against all 9 baselines. For a fair evaluation, we implement Flash attention for TokenGT as it was not originally supported. GraphGPS natively supports Flash attention, while Graphormer requires specific modifications to the attention matrix which are not currently supported.
(a) Training time for qm9.
(b) Allocated memory for qm9.
(c) Training time for mnist.
(d) Allocated memory for mnist.
@@ 4.5 Time and Memory Scaling
Figure 4: The elapsed time for training a single epoch for different datasets (in seconds), and the maximum allocated memory during this training epoch (GB). Different configurations are tested, while varying hidden dimensions and number of layers. qm9 has around 130K small molecules (a maximum of 29 nodes and 56 edges), dockstring has around 260K graphs that are around 6 times larger, and finally mnist has 70K graphs which are around 11-12 times larger than qm9. We use dummy integer features for TokenGT when benchmarking mnist. Graphormer runs out of memory for dockstring and mnist.
(a) Time-rank plot for all methods.
(b) Memory-rank plot for all methods.
Figure 5: All methods illustrated according to their achieved performance (rank) versus the total time spent training (minutes) in Panel (a) and the maximum allocated memory (GB) in Panel (b).
(a) The number of parameters (in millions) for all the methods and different configurations in terms of the hidden dimension and the number of layers.
(b) The time and memory utilisation for a selection of Graphormer configurations. The varying parameters are the number of layers, and the number of input graphs.
Figure 6: The number of parameters for all methods in Panel (a) and an illustration of the time and memory bottleneck in Graphormer in Panel (b). For Graphormer, even training on a single graph from hiv takes over 10 seconds and slightly under 5GB. Increasing the number of graphs to 8 and setting the batch size to 8 to ensure parallel computation increases the running time to around 80 seconds and slightly under 30GB. Extrapolating these numbers for a batch size of 8 to the entire dataset results in a training time of around 5 days for a single epoch.
@@
We report the training time for a single epoch and the maximum allocated memory during training for qm9 and mnist in Figure 4, and dockstring in Figure 1. In terms of training time, GCN and GIN are consistently the fastest due to their simplicity and low number of parameters (also see Figure 6(a)). ESA is usually the next fastest, followed by other graph transformers. The strong GNN baselines, particularly PNA, and to an extent GATv2 and GAT, are among the slowest methods. In terms of memory, GCN and GIN are again the most efficient, followed by GraphGPS and TokenGT. ESA is only slightly behind, with PNA, DropGIN, GATv2, and GAT all being more memory intensive, particularly DropGIN.
@@
We also order all methods according to their achieved performance and illustrate their rank relative to the total time spent training and the maximum memory allocated during training (Figure 5). ESA occupies the top left corner in the time plot, confirming its efficiency. The results are more spread out regarding memory, however the performance is still remarkable considering that ESA works over edges, whose number rapidly increases relative to the number of nodes that dictates the scaling of all other methods.
@@
Finally, we report the number of parameters for all methods and their configurations (Figure 6(a)). Apart from GCN, GIN, and DropGIN, ESA has the lowest number of parameters. GAT and GATv2 have rapidly increasing parameter counts, likely due to the concatenation that happens between attention heads, while PNA also increases significantly quicker than ESA. Lastly, we demonstrate and discuss one of the bottlenecks in Graphormer for the dataset hiv (Figure 6(b)). This lack of efficiency is likely caused by the edge features computed by Graphormer, which depend quadratically on the number of nodes in the graph, and on the maximum shortest path in the graph.
5 Discussion
@@
We presented an end-to-end attention-based approach for learning on graphs and demonstrated its effectiveness relative to tuned graph neural networks and recently proposed graph transformer architectures. The approach is free of expensive pre-processing steps and numerous other additional components designed to improve the inductive bias or expressive power (e.g., shortest paths, centrality, spatial and structural encodings, virtual nodes as readouts, expander graphs, transformations via interaction graphs, etc.). This shows huge potential and plenty of room for further fine-tuning and task-specific improvements. The interleaving operator inherent to ESA allows for vertical combination of masked and self-attention modules for learning effective token (i.e., edge or node) representations, leveraging relational information specified by input graphs while at the same time allowing to expand on this prior structure via self-attention. The recently released Flex Attention feature in PyTorch allows for further extensions via rich masking operators informed by graph structure while leveraging advancements in exact and efficient attention that have enabled our work.
@@
Our comprehensive evaluation shows that the proposed approach consistently outperforms strong message passing baselines and recently proposed transformer-based approaches for learning on graphs. The takeaway from our extensive study is that the proposed approach is well suited for being a simple and yet extremely effective starting point for learning on graphs. Moreover, the approach has favourable computational complexity and scales better than strong GNNs and some graph transformers. The approach also does well in transfer learning settings, possibly paving the way for more research on foundational models for drug discovery, where the problem arises frequently in property prediction for expensive high-fidelity experiments.

### v3

@@ Results
Table 6: The table reports the performance of tuned ESA+ models on a selection of six datasets, with different evaluation metrics to facilitate comparisons across published papers and leaderboards. The results for the full version of zinc (with 250,000 compounds) and a commonly used subset of zinc with 12,000 compounds are reported separately. Molecular models leverage random-walk structural encodings (RWSE). Table headers are presented in bold.
Model zinc zinc-full pept-str pept-fn mnist malnet
Metric MAE ($\downarrow$) MAE ($\downarrow$) MAE ($\downarrow$) AP ($\uparrow$) Accuracy ($\uparrow$) Accuracy ($\uparrow$)
ESA+ $0.05196916754671219$ $\pm$ $0.0009185022994316628$ $0.010894100678479495$ $\pm$ $0.00022839903839981288$ $0.2393$$\pm$ $0.0004$ $0.7357$$\pm$ $0.0036$ $98.917$$\pm$ $0.020$ $94.800$$\pm$ $0.424$
Node-level Benchmarks
@@
Node-level tasks are an interesting challenge for our proposed approach, as the PMA module is not needed and node-level propagation is the most natural strategy for learning node representations. To this end, we did prototype with an edge-to-node pooling module which would allow ESA to learn node embeddings, with good results. However, the approach does not currently scale to the millions of edges that some heterophilous graphs have (also see SI 6 and SI 7 for a discussion on extending ESA to millions of edges). For these reasons, we revert to the simpler node-set attention (NSA) formulation. Table 4 summarises our results, indicating that NSA performs well on all $11$ node-level benchmarks, including homophilous (citation), heterophilous, and shortest path tasks. We have adapted Graphormer and TokenGT for node classification as this functionality was not originally available, although they require integer node and edge features, restricting their use on some datasets (denoted by n/a in Table 4). NSA achieves the highest MCC score on homophilous and heterophilous graphs, but GCN has the highest accuracy on cora (Table 10). Some methods, for instance GraphSAGE [10.5555/3294771.3294869], are designed to perform particularly well under heterophily and to account for that we include GraphSAGE [10.5555/3294771.3294869] and Graph Transformer [ijcai2021p214] as the two top performing baselines from Platonov et al. [platonov2023a] in our extended results in Table 11. We also include Polynormer as another method dedicated to node classification. With a few exceptions, we outperform the baselines by a noticeable margin. Our results on chameleon stand out in particular, as well as on the shortest path benchmarks, where other graph transformers are unable to learn and even PNA fails to be competitive.
Effects of Varying the Layer Order and Type
@@
In Table 5, we summarise the results of an ablation relative to the interleaving operator, with different numbers of layers and layer orderings. Smaller datasets perform well with 4 to 6 feature extraction layers, while larger datasets with more complex graphs like mnist benefit from up to 10 layers. We have generally observed that the top configurations tend to include self-attention layers at the front, with masked attention layers in the middle and self-attention layers at the end, surrounding the PMA readout. Naive configurations such as all-masked layers or simply alternating masked and self-attention layers do not tend to be optimal for graph-level prediction tasks. This ablation experiment demonstrates the importance of vertically combining masked and self-attention layers for the performance of our model.
@@
Figure 3: A – D. The elapsed time for training all evaluated models for a single epoch and for different datasets (in seconds), and the maximum allocated memory during this training epoch (GB), while varying hidden dimensions and the number of layers. qm9 has around 130K small molecules (a maximum of 29 nodes and 56 edges), dockstring (Figure 1) has around 260K graphs that are around 6 times larger, and finally mnist has 70K graphs which are around 11-12 times larger than qm9. We use dummy integer features for TokenGT when benchmarking mnist. Graphormer runs out of memory for dockstring and mnist. E – F. All methods illustrated according to their achieved performance (rank) versus the total time spent training (E) and the maximum allocated memory (F). G. The number of parameters (in millions) for all the methods and different configurations in terms of the hidden dimension and the number of layers. H. The time and memory bottleneck in Graphormer. Even training on a single graph from hiv takes over 10 seconds and slightly under 5GB. Extrapolating these numbers for a batch size of 8 to the entire dataset results in a training time of 5 days for a single epoch.
Enhancing ESA for State-of-the-Art Performance
So far, we have focused on establishing that ESA is a strong foundation for graph learning and have not leveraged positional and structural encodings or other helper mechanisms (with the exception of zinc in Table 1). These models were already outperforming or matching the state-of-the-art on several datasets. To explore the potential and compatibility of ESA with additional transformations and components, we further tune hyperparameters such as the optimiser and explore the utility of structural encodings [grötschla2024benchmarkingpositionalencodingsgnns] and/or problem-specific modifications of input graphs (i.e. explicit incorporation of ring membership as node features for zinc [menegaux2023selfattention]). For clarity, we refer to this extension as ESA+ and report our results in Table 6. The results on all six datasets are state-of-the-art and top the online leaderboards at the time of writing (March 2025).
Time and Memory Scaling
The technical aspects of ESA are discussed in Methods, Time and Memory Scaling, where we also cover deep learning library limitations and possible optimisations (also see SI 6 and SI 7). Although efficient attention implementations have linear memory complexity, storing the edge adjacency matrices in a dense format is currently a bottleneck imposed by available frameworks. Here, we have empirically evaluated the time and memory scaling of ESA against all 9 baselines. For a fair evaluation, we implemented Flash attention for TokenGT as it was not originally supported. GraphGPS natively supports Flash attention, while Graphormer requires specific modifications to the attention matrix that are not currently supported.
@@
We report the training time for a single epoch and the maximum allocated memory during training for qm9 and mnist in Figure 3A to D, and dockstring in Figure 1. In terms of training time, GCN and GIN are consistently the fastest due to their simplicity and low number of parameters (also see Figure 3G). ESA is usually the next fastest, followed by other graph transformers. The strong GNN baselines, particularly PNA, and to an extent GATv2 and GAT, are among the slowest methods. In terms of memory, GCN and GIN are again the most efficient, followed by GraphGPS and TokenGT. ESA is only slightly behind, with PNA, DropGIN, GATv2, and GAT all being more memory intensive, particularly DropGIN.
@@
We also order all methods according to their achieved performance and illustrate their rank relative to the total training time and the maximum memory allocated during training (Figure 3E and F). ESA occupies the top left corner in the time plot, confirming its efficiency. The results are more spread out regarding memory; however, the performance is still remarkable considering that ESA works over edges, whose number rapidly increases relative to the number of nodes that dictates the scaling of all other methods.
@@
Finally, we investigate and report the number of parameters for all methods and their configurations (Figure 3G). Apart from GCN, GIN, and DropGIN, ESA has the lowest number of parameters. GAT and GATv2 have rapidly increasing parameter counts, likely due to the concatenation that occurs between attention heads, while PNA also increases significantly quicker than ESA. Lastly, we demonstrate and discuss one of the bottlenecks in Graphormer for the dataset hiv (Figure 3H). This lack of efficiency is likely caused by the edge features computed by Graphormer, which depend quadratically on the number of nodes in the graph and on the maximum shortest path in the graph.
Explainability
Explainability is one aspect where attention models excel compared to other architectures, and since ESA is fundamentally edge-based, analysing attention scores might lead to different insights compared to traditional models. Here, we investigate two ways of interpreting the learnt attention weights/scores of ESA and visualise the results based on the qm9 model from Table 1, using the following architecture: MMMMMMSPS. This choice is motivated by the availability of several quantum properties with significantly different physical characteristics, and in particular (1) HOMO (highest occupied molecular orbital) energy, an intensive and localised property, and (2) u0 (internal energy at 0K), an extensive and global property. As the HOMO can be highly localised in a region of the molecule, while the internal energy should be evenly distributed across all atoms, we expect this dynamic to be reflected in the weights learnt by the model. More specifically, in the case of HOMO we hypothesise that only a few attention scores would dominate for each molecule, while for u0 the scores should be dispersed relatively uniformly over all the atoms.
To quantify this notion, we leveraged the Gini coefficient, a measure of inequality that ranges from 0 to 1. A value of 0 indicates equality among all values, while a value of 1 indicates maximal inequality, where a single value dominates and the rest are 0. Our results are illustrated in Figure 4 for all 7 encoder layers and show that while the attention scores start from an even spread for both HOMO and u0, in the case of HOMO we notice progressively more peaked and dominant scores until the last layer, while for u0 the opposite is true, with the latter layers exhibiting the most dispersed attention.
For our second analysis, we investigated whether the bonds corresponding to the top attention scores of the HOMO model contribute to the ground truth molecular orbitals calculated from the 3D conformations available in the qm9 dataset. Our protocol and visualisations are covered in SI 16, showing that the model indeed learns physically relevant relationships for a variety of molecules.
Discussion
@@
We presented an end-to-end attention-based approach for learning on graphs and demonstrated its effectiveness relative to tuned graph neural networks and recently proposed graph transformer architectures. The approach is free of expensive pre-processing steps and numerous other additional components designed to improve the inductive bias or expressive power (e.g. shortest paths, centrality, spatial and structural encodings, virtual nodes as readouts, expander graphs, transformations via interaction graphs, etc.). This shows huge potential and plenty of room for further fine-tuning and task-specific improvements. The interleaving operator inherent to ESA allows for vertical combination of masked and self-attention modules for learning effective token (i.e. edge or node) representations, leveraging relational information specified by input graphs while at the same time allowing to expand on this prior structure via self-attention.
Figure 4: A – B. Gini coefficients for attention scores on two quantum properties: HOMO (highest occupied molecular orbital energy) in A and u0 (internal energy at 0K) in B. The attention scores for each layer are first averaged over the number of heads.
ESA utilises line graphs [whitney32] where the edges from an input graph are transformed into nodes in the line graph, establishing connections if they share a common node in the original graph. This transformation preserves isomorphisms, meaning that if two graphs are isomorphic, their line graphs are as well; however, the reverse is not necessarily true. Notably, some non-isomorphic graphs that are indistinguishable by the $1$-WL test become distinguishable through their line graph embeddings (see SI 17 and 4), possibly highlighting richer information conveyed by edge tokens as opposed to node tokens in some tasks. Line graphs naturally emphasise important motifs [whitney32, Harary1960SomePO, bauer19821] like triangles, cycles, and cliques, which are critical for tasks such as predicting molecular properties. Furthermore, line graphs tend to balance vertex degree variance, maintaining a more uniform distribution which enhances the regularity of message passing. This edge-focused communication strategy aligns with previous explorations in GNNs, such as in Chemprop [chemprop], which hypothesised improved accuracy resulting from handling information propagation more effectively than node-centric message passing.
@@
Our comprehensive evaluation shows that the proposed approach consistently outperforms strong message passing baselines and recently proposed transformer-based approaches for learning on graphs. The takeaway from our extensive study is that the proposed approach is well suited to be a simple and yet extremely effective starting point for learning on graphs. Indeed, preliminary results on enhancing ESA with positional encodings indicate state-of-the-art results on several datasets. Moreover, the approach has favourable computational complexity and scales better than strong GNNs and some graph transformers. ESA also does well in transfer learning settings, possibly paving the way for more research on foundational models for structured data and drug discovery in particular, where the problem arises frequently in property prediction for expensive high-fidelity experiments.
The widespread approach of storing attention masks as dense tensors is likely caused by the limited interest in structured masking, since the main driving force behind efficient attention implementations is scaling up large language models (LLMs), where the only masking requirements are simple causal masks or similar pre-defined patterns. In principle, one only needs to load and apply the masking information per chunk, and in practice this manipulation would lead only to a small overhead while avoiding hundreds of millions of computations. Such an implementation is not trivial, as it requires knowledge of GPU architectures and low-level programming abstractions like Triton [10.1145/3315508.3329973] and CUDA [10.1145/1401132.1401152]. As such, it is not within the scope of our current work. The very recent Flex Attention project [dong2024flexattentionprogrammingmodel] is partially motivated by our initiative and aims to extend and streamline efficient attention operations, including sparse patterns. However, such prototypes are still not fully functional. Nevertheless, even our current implementation covers a wide diversity of application domains, such as computational chemistry (small molecules and peptides), computer vision, bioinformatics, social networks, and many others.
Methods
In this section, we provide a high-level overview of our architecture and describe the masked attention module that allows information to propagate across edges as primitives. An alternative implementation for the more traditional node-based propagation is also presented. The section starts with a formal definition of the masked self-attention mechanism and then proceeds to describe our end-to-end attention-based approach to learning on graphs, involving the encoder and the pooling components (illustrated in Figure 2).
Masked Attention Modules
We first introduce the basic notation and then give a formal description of the masked attention mechanism with a focus on the connectivity pattern specific to graphs. Following this, we provide algorithms for an efficient implementation of masking operators using native tensor operations.
As in most graph learning settings, a graph is a tuple $\mathcal{G}=(\mathcal{N},\mathcal{E})$ where $\mathcal{N}$ represents the set of nodes (vertices), $\mathcal{E}\subseteq\mathcal{N}\times\mathcal{N}$ is the set of edges, and $N_{n}=|\mathcal{N}|$, $N_{e}=|\mathcal{E}|$. The nodes are associated with the feature vectors $\mathbf{n}_{i}$ of dimension $d_{n}$ for all nodes $i\in\mathcal{N}$, and $d_{e}$-dimensional edge features $\mathbf{e}_{ij}$ for all edges $(i,j)\in\mathcal{E}$. The node features are collected as rows in a matrix $\mathbf{N}\in\mathbb{R}^{N_{n}\times d_{n}}$, and similarly for edge features in $\mathbf{E}\in\mathbb{R}^{N_{e}\times d_{e}}$. The graph connectivity can be specified as an adjacency matrix $\mathbf{A}$, where $\mathbf{A}_{ij}=1$ if $(i,j)\in\mathcal{E}$ and $\mathbf{A}_{ij}=0$ otherwise. The edge list (edge index) representation is equivalent but more common in practice.
where $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ are projections of the edge representations to queries, keys, and values, respectively. This is a minor extension of the standard scaled dot product attention via the additive mask $\mathbf{M}$ that can censor any of the pairwise attention scores, and $d_{k}$ is the key dimension. The generalisation to masked multi-head attention follows the same steps as in the original transformer [NIPS2017_3f5ee243]. In the following sections, we refer to this function as $\text{MultiHead}(\mathbf{Q},\mathbf{K},\mathbf{V},\mathbf{M})$.
The masked self-attention for graphs can be seen as graph-structured attention, and an instance of a generalised attention mechanism with custom attention patterns [shirzad2023exphormer]. More specifically, in ESA the attention pattern is given by an edge adjacency matrix rather than allowing for interactions between all set items. Crucially, the edge adjacency matrix can be efficiently computed both for a single graph and for batched graphs using exclusively tensor operations. The case for a single graph is covered in the left-hand side of Introduction through the edge_adjacency function. The first 3 lines of the function correspond to getting the number of edges, then separating the source and target nodes from the edge adjacency list (also called edge index), which is equivalent to the standard adjacency matrix of a graph. The source and target node tensors each have a dimension equal to the number of edges ($N_{e}$). Lines 7-9 add an additional dimension and efficiently repeat (expand) the existing tensors to shape $(N_{e},N_{e})$ without allocating new memory. Using the transpose, line 11 checks if the source nodes of any two edges are the same, and the same for target nodes on line 12. On line 13, cross connections are checked, where the source node of an edge is the target node of another, and vice versa. The operations of lines 11-14 result in boolean matrices of shape $(N_{e},N_{e})$ which are summed for the final returned edge adjacency matrix. The right-hand side panel of Introduction shows the case where the input graph represents a batch of smaller graphs. This requires an additional batch mapping tensor that maps each node in the batched graph to its original graph, and carefully manipulating the indices to create the final edge adjacency mask of shape $(B,L,L)$, where $L$ is the maximum number of edges in the batch.
Since in ESA attention is computed over edges, we chose to separate source and target node features for each edge, similarly to lines 4-5 of Introduction, and concatenate them to the edge features:
Having defined the mask generation process and the masked multi-head attention function, we next define the modular blocks of ESA, starting with the Masked Self Attention Block (MAB):
$\mathbf{\overline{X}}$ $=\text{LayerNorm}(\mathbf{X})$ (5)
where the mask $\mathbf{M}$ is computed as in Introduction and has shape $B\times L\times L$, with $B$ the batch size, and MLP is a multi-layer perceptron. A Self Attention Block (SAB) can be formally defined as:
$\text{SAB}(\mathbf{X})=\text{MAB}(\mathbf{X},\mathbf{0})$ (6)
ESA Architecture
Our architecture consists of two components: i) an encoder that interleaves masked and self-attention blocks to learn an effective representation of edges, and ii) a pooling block based on multi-head attention, inspired by the decoder component from the set transformer architecture [lee2019set]. The latter comes naturally when one considers graphs as sets of edges (or nodes), as is done here. ESA is a purely attention-based architecture that leverages the scaled dot product mechanism proposed by Vaswani et al. [NIPS2017_3f5ee243] through masked and standard self-attention blocks (MABs and SABs) and a pooling by multi-head attention (PMA) block.
Here, AB refers to an attention block that can be instantiated as an MAB or SAB.
where $\mathbf{S}_{k}$ is a tensor of $k$ learnable seed vectors that are randomly initialised and $\text{SAB}^{p}(\cdot)$ is the application of $p$ SABs. Technically, it suffices to set $k=1$ to output a single representation for the entire graph. However, we have empirically found it beneficial to set it to a small value, such as $k=32$. Moreover, this change allows self-attention (SABs) to further process the $k$ resulting representations, which can be simply summed or averaged due to the small $k$. Contrary to classical readouts that aggregate directly over set items (i.e. nodes), pooling by multi-head attention performs the final aggregation over the embeddings of learnt seed vectors $S_{k}$. Whereas tasks involving node-level predictions require only the Encoder component, predictive tasks involving graph-level representations require all the modules, both the encoder and pooling by multi-head attention. The architecture in the latter setting is formally given by
$\mathbf{Z}_{\text{out}}=\text{PMA}_{k,p}(\text{Encoder}(\mathbf{X},\mathbf{M})+\mathbf{X})$ (10)
As optimal configurations are task specific, we do not explicitly fix all the architectural details. For example, it is possible to select between layer and batch normalisation, a pre-LN or post-LN architecture [10.5555/3524938.3525913], or standard and gated MLPs, along with GLU variants, e.g. SwiGLU [DBLP:journals/corr/abs-2002-05202].
Time and Memory Scaling
ESA is enabled by recent advances in efficient and exact attention, such as memory-efficient attention and Flash attention [DBLP:journals/corr/abs-2112-05682, dao2022flashattention, dao2024flashattention, shah2024flashattention]. Memory-efficient implementations with arbitrary masking capabilities exist in both PyTorch and xFormers [NEURIPS2019_9015, xFormers2022]. Efficient attention avoids materialising the $N\times N$ square attention matrix, instead working over smaller chunks and saving partial aggregates that can recover the actual attention scores. Thanks to chunking and several optimisations, exact attention can be computed with linear memory requirements. Furthermore, although in theory the time complexity is still given by a quadratic number of computations, in practice GPU optimisations such as chunking, fused kernels, reduced memory bottlenecks (better bandwidth utilisation), and others can lead to forward passes that are up to ×$\times 25$ times faster, with comparable uplifts for the backward pass [dao2024flashattention, shah2024flashattention]. The complexity of ESA directly corresponds to that of the underlying attention implementation, since that is the main learning mechanism, augmented with relatively cheap MLPs and normalisation layers. A precise complexity analysis would have to include specific GPU architecture details that change from one GPU type to another, as well as a per-case discussion based on the specific chunk size, sequence length (here, number of edges), total hidden dimension, and number of heads. This is because the kernels are available only for certain combinations of sequence lengths, dimensions, and heads, and they differ in performance.
All attention operations in ESA except the cross-attention in PMA are based on self-attention, and our method directly benefits from all of these advances. Moreover, our cross-attention is performed between the full set of size $L$ and a small set of $k\leq 32$ learnable seeds, such that the complexity of the operation is $O(kL)$ even for standard attention. Indeed, currently the main bottleneck is not the attention computation itself, but storing the dense edge adjacency masks. Unfortunately, all the available libraries that support masking and efficient attention kernels require a dense matrix mask, which demands memory quadratic in the number of edges. An evident, but not yet available, optimisation could amount to storing the masks in a sparse tensor format. To visualise the benefits of this optimisation, we estimate the cost of storing the sparse edge adjacency information in addition to end-to-end training a model with a full self-attention layer (no masking). The performance of this hypothetical model is visualised in SI 6 and shows that ESA could scale to graphs of 2.6 million edges on a single GPU with 40GB memory, even without masking hundreds of millions of unnecessary calculations. Despite some of the current limitations, we have successfully trained ESA models with batch sizes of 8 graphs with approximately 30,000 edges each (e.g. dd in Table 3), roughly equivalent to a single graph with 240,000 edges. Further limitations and possible optimisations are discussed in SI 7.
Data Availability
All the datasets used throughout the paper are publicly available through different hosting services, as indicated in the main text and supplementary materials. The majority of datasets are accessed via PyTorch Geometric. Other sources include DOCKSTRING [GarciaOrtegon2022], heterophily datasets, accurate GW frontier orbital energies of 134 kilo molecules of the QM9 dataset [Fediai2023data], Open Catalyst Project. All instances of synthetic data are reproducibly generated by the code available in our GitHub repository; processed datasets (e.g. infected Erdős–Rényi graphs) are also available in our repository.
Code Availability
The source code that enables all experiments to be reproduced, as well as code documentation, examples, and set-up instructions are available in our GitHub repository: https://github.com/davidbuterez/edge-set-attention.